OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 7 - 9x - (x + 12) less than or greater to 25? x greater than or equal to -3 x less than or greater to -3 -3 less than or greater to x A number line with a closed circle on negative 3 and shading to the right. @Preetha @Luis_Rivera @hba @waterineyes

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you sure that the inequality sign is "less than or greater to"

OpenStudy (anonymous):

Check the statement of your question once again..

OpenStudy (anonymous):

@waterineyes @jim_thompson5910 it's a typo. It's supposed to be 'less than or equal to'. Sorry.

jimthompson5910 (jim_thompson5910):

that's ok, thanks for clearing it up

jimthompson5910 (jim_thompson5910):

7 - 9x - (x + 12) <= 25 7 - 9x - x - 12 <= 25 -10x - 5 <= 25 -10x <= 25 + 5 -10x <= 30 I'll let you finish

OpenStudy (anonymous):

Just open up the brackets on left and solve it futher: \[7 - 9x - x - 12 \le 25 \implies -10x - 5 \le 25 \implies -10x \le 30\]

OpenStudy (anonymous):

Here: when multiplying by "-", you will have to take extra care..

OpenStudy (anonymous):

Like this? http://www.sketchtoy.com/29325555

jimthompson5910 (jim_thompson5910):

good, the sign will flip because you're dividing both sides by a negative number

OpenStudy (anonymous):

Thanks guys!